OpenStudy (anonymous):

M is the midpoint of JK, find the coordinates of K. J(-8,4), M(-1,1)

OpenStudy (anonymous):

\[\large midpoint (x,y) = (\frac{x_1 + x_2}{2}, \frac{y_1 + y_1}{2})\]

OpenStudy (anonymous):

Here you are given with Midpoint*x,y) = (-1, 1)

OpenStudy (anonymous):

So, \[x = \frac{x_1 + x_1}{2} \implies -1 = \frac{-8 + x_2}{2}\] Find \(x_2\) here..